OpenStudy (dellgato):

Find the perimeter of a triangle with side lengths (6x+2) in, (x-12) in, and (3x+15) in.

Nnesha (nnesha):

perimeter of a triangle  \[\large \rm P=x+y+z\]

OpenStudy (dellgato):

How do ! add them together though

Nnesha (nnesha):

combine like terms add the coefficients of identical variables

Nnesha (nnesha):

for example 2x+7x = (2+7)x =9x

OpenStudy (dellgato):

Ok so after adding the like terms I get 10x and 5

Nnesha (nnesha):

10x and 5 is it +5 or -5 ?

OpenStudy (dellgato):

+5?

Nnesha (nnesha):

yes right good!

Nnesha (nnesha):

don't forget the unit !

OpenStudy (dellgato):

Ok so how do i right the answer (10x+5) or (10x-5)

OpenStudy (dellgato):

I write the answer* excuse me

Nnesha (nnesha):

did you get -5 or +5 ?? :=))

OpenStudy (dellgato):

OHHHHHH ok thanks a lot

Nnesha (nnesha):

\[\huge\rm P=(6x+2)in+(x-12)in+(3x+15)in\] there are + signs so you can remove the parentheses \[\large\rm P=6x+2~in+x-12~in+3x+15~in\] combine like terms

Nnesha (nnesha):

so how would you write your final answer ?

OpenStudy (dellgato):

(10x+5)

OpenStudy (dellgato):

Am I correct

Nnesha (nnesha):

\[\large\rm P=(6x +2 )\color{red}{in}+(x-12)\color{Red}{in} +(3x+15)\color{red}{in}\] don't forget the unit.. \[\large\rm P=(6x +2 +x-12+3x+15) \color{Red}{in}\]

Nnesha (nnesha):

yes that's correct 10x+5 `in`

OpenStudy (dellgato):

10x+5in

OpenStudy (dellgato):

Thanks

Nnesha (nnesha):

ye or you can write it as (10x+5)in

Nnesha (nnesha):

for example if the sides are 4 m , 5m , 6m then the perimeter would be P=4m+5m+6m=(4+5+6)m = 15m the unit is important it workz the same way like `like terms`

Nnesha (nnesha):

yw